OpenStudy (anonymous):

Help with 2nd page please? MEDAL REWARDED!!

OpenStudy (anonymous):

@Mertsj ?

OpenStudy (anonymous):

I am not sure what to do?

OpenStudy (mertsj):

Essentially you have two choices for each table. The first choice is: square x and multiply it by a number to get y.

OpenStudy (anonymous):

Okay

OpenStudy (mertsj):

Let's try that for the first problem. (-2,8) is the first ordered pair. If we square x we get 4. If we multiply 4 by 2 we get 8.

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

Do we need to make a table?

OpenStudy (anonymous):

Or something similar?

OpenStudy (mertsj):

Let's see if that works for the second ordered pair (-1,5) If we square x we get 1. If we multiply that by 4 we get 4. OH NO!!! y is not 4. So that does not work for the first table.

OpenStudy (anonymous):

So it is not linear?

OpenStudy (mertsj):

So it is NOT y = ax^2. Therefore it must be y = x^2+c

OpenStudy (anonymous):

Oh okay

OpenStudy (mertsj):

These are parabolas, sweetheart. Not lines.

OpenStudy (anonymous):

Oh sorry

OpenStudy (mertsj):

That's why it says "parabo-lems"

OpenStudy (anonymous):

Lol

OpenStudy (mertsj):

Can you figure out the second one?

OpenStudy (anonymous):

So what do we put for the first problem then?

OpenStudy (anonymous):

Our teacher told us to write a verbal description and an equation.

OpenStudy (anonymous):

Based on the parabolas shown.

OpenStudy (mertsj):

If you have to write all that I would put : \[(-2)^2+4=8\] \[(-1)^2+4=5\] \[(0)^2+4=4\] \[(1)^2+4=5\] \[(2)^2+4=8\] So this parabola is of the form : \[y=x^2+c \]; and c = 4

OpenStudy (anonymous):

Thanks for the steps! :D

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

So it is y=x^2+4 Or y=4x^2? @Mertsj

OpenStudy (mertsj):

Did we square x and multiply by 4 or square x and add 4?

OpenStudy (anonymous):

Add 4?

OpenStudy (mertsj):

Yes. If you bother to read my extensive post above, you will see that each one says +4

OpenStudy (anonymous):

Yes, I did bother too, lol

OpenStudy (anonymous):

Thanks!

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

Help with the second one, second question please?

OpenStudy (mertsj):

Square each x and see what you have to do to that answer to get y

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

Or each coordinate right?

OpenStudy (anonymous):

Lol

OpenStudy (mertsj):

each x coordinate...the first number in the ordered pair.

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

Can you help me step by step, please, because I have a test tomorrow and I need some practice.

OpenStudy (anonymous):

Oh for the problem you just solve is wrong. I meant the second page not the first page.

OpenStudy (mertsj):